# Toward A Causal Framework for Modeling Perception

### v3

@@ 6.2 Loaded Attributes
Similarly, in previous sections we did not discuss in detail sensitive attributes, like gender or race, and their role in perception. We argue that these attributes are prone to induce perception, as they are summaries of complex historical and social processes (Bonilla-Silva 1997; Sen and Wasow 2016). As such, they are likely to affect how individuals are perceived by others. We are referring, e.g., to the conceptual difference between describing an individual as female versus feminine (Hu and Kohler-Hausmann 2020), both of which are based on the attribute gender. Female refers to a category of gender while feminine refers to a set of behavioral expectations attributed to females: i.e., the biological or phenotypic label versus the socially constructed identity.
We describe these attributes as loaded because they are likely to lead to different interpretations among receivers. If $X$ is a loaded attribute, then it should be easier for a receiver $R$ to evoke individual experience about $X$. Broadly, we view loaded attributes as attributes that thrive on stereotypes of social categories shared and maintained by the receivers. A social category is the result of classifying people into groups over shared perceived identities (Bowker and Star 1999). We refer to a social category as a social construct when the classification is also used purposely to enforce exclusionary policies (Mallon 2007). Sensitive attributes are examples of social constructs. A stereotype refers to the cognitive representation people develop about a particular social category, based around beliefs and expectations about probable behaviors, features and traits, which can translate into implicit or explicit attitudes that materialize into bias (Beukeboom and Burgers 2019; Johnson 2020).
The difficulty, as discussed in Section 5.3, is how we model such distinction causally and, further, how it translates into potential disagreement among decision makers. The causal perception framework offers a way to formalize these overloaded attributes as elements of individual experience from the perspective of the receiver: SCMs condition how receivers interpret information, which is particularly important given the potential biases they may hold toward certain social groups.
@@
As implied in Examples 6 and 7, the officer consideration of the address, $Z$, of each applicant serves as an attempt to infer their socioeconomic background. By associating $Z$ with $X_{1}$ and $X_{2}$ and drawing on stereotypes (e.g., correlations between wealth and SAT scores), an officer may develop biased inferences, which ultimately lead to favor one applicant over the other.
@@ 6.3 The Human-in-the-Loop
The following example, while centered on automated decision-making (ADM), illustrates the same interpretative challenges discussed in the previous examples for scenarios (i) and (ii). In this context, we can view the admissions officer as one receiver and her supervisor as another. The natural question to ask then is “what happens if they disagree on the $k$ applicants selected by $f$?” Such situations highlight a broader issue in ML pipelines: decision processes often culminate in human oversight. From a fairness perspective, it is essential to develop frameworks that can accommodate subjective interpretations and context-sensitive notions of fairness (Srivastava, Heidari, and Krause 2019; Yaghini, Krause, and Heidari 2021; Jung et al. 2021).
@@
Consider scenario (iii) in Example 1. Given the top-$k$ applicants by the ML model $f$, suppose that the admissions officer is asked by the college to account for the chosen applicants by $f$. The officer uses xAI tools, like feature importance methods, to understand the model’s decisions and explain them to a supervisor. Assume that the supervisor also has access to $f$ and the xAI tools.
As studied by Bertrand et al. (2022), suppose that in Example 10 the officer and her supervisor interpret differently the feature importance explanations. This divergence highlights a key challenge in xAI: how to design explanations that minimize interpretive ambiguity among users. Similar to nudging in behavioral economics (Gigerenzer 2018), which was inspired by Tversky and Kahneman’s work on cognitive biases (Thaler and Sunstein 2008), future xAI tools should aim to anticipate a range of plausible user interpretations and proactively guide users toward a shared, intended understanding. The proposed frameworks allows to formalize such set of user-based interpretations in a that is compatible with ML applications, e.g., with xAI tools.
@@ 7 Conclusion
In this work, we defined perception of probabilities and of implied probabilities. We treated perception, broadly, as a disagreement between individual probabilistic interpretations, under the premise that human decision-making is inherently probabilistic. For the latter, which we term causal perception, we used SCM Pearl (2009) to formalize disagreement between observed and interventional (read, hypothetical) distributions, allowing us to capture the range of causal probabilistic reasoning. Further, we defined two kinds of causal perception based on when the individuals disagree on the causal graph (structural perception) and when the individuals disagree on the causal effect given the same casual graph (parametrical perception). Additionally, we proposed an initial approach to how individuals construct these SCMs based on their experiences. Finally, we outlined possible fairness research areas that could benefit from the proposed framework.
Causal perception is useful in contexts where multiple interpretations (as well as representations) of information can occur. It is, in turn, useful for fairness problems involving multiple experts that interact with ML systems.
The work of this paper is conceptual in nature. Future work should aim to extend the formalization beyond the limitations of causality, and validate the framework empirically through realistic case studies.
@@ Appendix A Ethical Statement
We did not face ethical challenges when drafting the paper.
@@ Appendix B Acknowledgments
We are greatly thankful to Riccardo Massida for his discussions, insights, and suggestions regarding the idea of Causal Perception. We also would like to thank Mayra Russo, Maria-Esther Vidal, Laura State, and Andrea Pugnana for their feedback on earlier versions. Work supported by the European Union (EU)’s Horizon Europe research and innovation program for the project FINDHR (g.a. 101070212), and under the Horizon 2020 Marie Sklodowska-Curie Actions research and innovation program for the project NoBIAS (g.a. 860630). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the EU. Neither the EU nor the granting authority can be held responsible for them.

### v4

@@ 6.2 Loaded Attributes
Similarly, in previous sections we did not discuss sensitive attributes, like gender or race, and their role in perception. We argue that these attributes are prone to induce perception, as they are summaries of complex historical and social processes (Bonilla-Silva 1997; Sen and Wasow 2016). As such, they are likely to affect how individuals are perceived by others. We are referring, for example, to the conceptual difference between describing an individual as female versus feminine (Hu and Kohler-Hausmann 2020), both of which are based on the attribute gender. Female refers to a category of gender while feminine refers to a set of behavioral expectations attributed to females: i.e., the biological or phenotypic label versus the socially constructed identity. We describe these attributes as loaded because they are likely to lead to different interpretations among receivers. If $X\in\mathbf{X}$ is a loaded attribute, then it should be easier for a receiver $R$ to evoke individual experience about $X$.
We regard loaded attributes as attributes that thrive on stereotypes of social categories shared and maintained by the receivers. A social category is the result of classifying people into groups over shared perceived identities (Bowker and Star 1999). We refer to a social category as a social construct when the classification is also used purposely to enforce exclusionary policies (Mallon 2007). Sensitive attributes are examples of social constructs. A stereotype refers to the cognitive representation people develop about a particular social category, based around beliefs and expectations about probable behaviors, features and traits, which translate into implicit or explicit attitudes that materialize into bias (Beukeboom and Burgers 2019; Johnson 2020).
The main difficulty, as discussed in Section 5.3, is how to model causally such distinction and, further, how it translates into a potential disagreement among the decision makers. The causal perception framework offers a way to formalize these overloaded attributes as elements of individual experience from the perspective of the receiver. The SCMs condition how receivers interpret information, which is particularly important given the potential biases they may hold toward certain social groups.
@@
As implied in Examples 6 and 7, the admissions officers’ consideration of the zip code, $Z$, of each applicant serves as an attempt to infer their socioeconomic background. By associating $Z$ with $X_{1}$ and $X_{2}$ and drawing on stereotypes (like correlations between wealth and SAT scores), an officer may develop biased inferences, which ultimately lead to favor one applicant over the other.
@@ 6.3 The Human-in-the-Loop
The next example, while centered on automated decision-making (ADM), illustrates the same interpretative challenges discussed in the previous ones for scenarios (i) and (ii). We view the admissions officer as one receiver and her supervisor as another receiver. A natural question to ask then is “What happens if they disagree on the $k$ applicants selected by ff?” in scenario (iii). Such situation highlights a broader issue with ML pipelines: the final decision often culminates in human oversight. From a fairness perspective, it is essential to develop frameworks that accommodate subjective interpretations and context-aware notions of fairness (Srivastava, Heidari, and Krause 2019; Yaghini, Krause, and Heidari 2021; Jung et al. 2021).
@@
Consider scenario (iii) in Example 1. Given the top-$k$ applicants by the ML model $f$, suppose that the admissions officer is asked by the college to explain for the $k$ chosen applicants by $f$. The officer uses xAI tools, like feature importance methods, to understand the model’s outcome and explain it to her supervisor. Assume that the supervisor also has access to $f$ and the same xAI tools. Suppose the officer and her supervisor interpret differently the feature importance explanations.
Example 10 is based on studies such as Bertrand et al. (2022) that consider the limitations of xAI methods under the threat of human cognitive biases. This divergence between officer and supervisor highlights a key challenge in xAI: how can we design explanations that minimize interpretive ambiguity among users? Similar to nudging in behavioral economics (Gigerenzer 2018), which was inspired by Tversky and Kahneman’s work on cognitive biases (Thaler and Sunstein 2008), future xAI tools should aim to anticipate a range of plausible user interpretations and proactively guide users toward a shared, intended understanding. The causal perception framework allows to formalize such set of user-based interpretations in a way that is compatible with ML applications and xAI tools.
@@ 7 Conclusion
In this work, we treated perception as a disagreement between individual probabilistic interpretations of the same bit of information, and defined perception of probabilities and perception of implied probabilities. For the latter, which we termed causal perception, we used structural causal models (SCMs) to formalize disagreement between observed and interventional (read, hypothetical) distributions. Further, we formalized two kinds of causal perception based on when individuals disagree on the causal graph (structural perception) and when they disagree on the causal effect(s) given the same casual graph (parametrical perception). Furthermore, we proposed a first approach to how individuals construct these disagreeing individual SCMs through the processes of categorization and signification.
We also outlined fairness research areas that benefit from the proposed framework. Notably, causal perception is useful in contexts in which multiple interpretations and representations of information occur by experts. It is, in turn, useful for problems involving diverse stakeholders that interact with the ML system. The framework allows to formally position fairness in such a setting as well as to reformulate sensitive attributes as loaded with perceptions.
This paper is, above all, conceptual. Next steps include, for instance, using ontologies (Gruber 1995) and relational learning (Salimi et al. 2020)—with Kahneman and Miller (1986)’s norm theory in mind—to construct in practice the SCMs that denote individual experience. Revisiting causal perception under Pearl’s three levels of reasoning (here, we simplified it to two levels) is also a next step (Bareinboim et al. 2022). In general, future work should validate the framework empirically through case studies. Additionally, causal perception assumes an explicit level of causal reasoning among individuals that may not be realistic as well as assumes SCMs as individual prior knowledge. Future work should evaluate how practical and scalable these individual SCMs are for the context in question.
@@ Appendix A Ethical Statement
We did not face ethical challenges when drafting this paper.
@@ Appendix B Acknowledgments
We are indebted to Riccardo Massida for his discussions, insights, and suggestions. We would like to thank also Mayra Russo, Maria-Esther Vidal, and Andrea Pugnana for their feedback in earlier versions of this paper. Salvatore Ruggieri is supported by the European Union (EU)’s Horizon Europe research and innovation program for the project FINDHR (g.a. 101070212). Views and opinions expressed are those of the authors only and do not necessarily reflect those of the EU. Neither the EU nor the granting authority can be held responsible for them.